(3) (i) A broken pipe at an oil rig off the east coast of Trinidad produces a circular oil slick that is S meters thick at a distance x meters from the break. It difficult to measure the thickness of the slick directly at the source owing to excess turbulence, but for x >0 they know that s(x) = x^2/2 + 5/3x/x^3 + x^2 + 2x If the oil slick is assumed to be continuously distributed, how thick is expected to be at the source? (ii) If f(x) = {4x + 7 1< x < 2, 4x^2 - 1 2< x < 4 , determine whether the function f (x) is continuous throughout its domain? 


1
Expert's answer
2021-11-14T16:45:19-0500

I.

"\lim\limits_{x\to0} \space \frac{\frac{x^2}{2}+\frac{5}{3}x}{x^3+x^2+2x}\\

\lim\limits_{x\to0} \space \frac{X(\frac{x}{2}+\frac{5}{3})}{X(x^2+x+2}\\

\lim\limits_{x\to0} \space \frac{\frac{x}{2}+\frac{5}{3}}{x^2+x+2}"

"\frac{\frac{x}{2}+\frac{5}{3}}{0^2+0+2}=\frac{5}{6}"

Oil slick is "\frac{5}{6}"m thick of source.


ii.

"f(x)=\begin{bmatrix}
 4x+7 & 1
≤
x


≤2\\
 4x^2-1& 2<x≤4
\end{bmatrix}"


"\lim\limits_{x\to2^-}(4x+7)\\=4(2)+7\\=13"

"\lim\limits_{x\to2^+}(4x^2-1)\\=4(2)^2-1\\=15"

"\lim\limits_{x\to2^-} f(x)\neq \lim\limits_{x\to2^+} f(x)"

f(x) is discontinuous at x=2.